OpenStudy (anonymous):

Karen is eleven years less than four times John’s age. If Karen is 49, how old is John?

OpenStudy (cgreenwade2000):

We have everything we need to solve this. Just set up an equation. K = 4J -11 K=49 Solve.

OpenStudy (anonymous):

K=4j-11 49=4j-11 -11 -11 ----------- 38=4j -- -- 4 4 j=9.5

OpenStudy (cgreenwade2000):

No sam. You would add 11 on both sides.

OpenStudy (anonymous):

oh yea

OpenStudy (anonymous):

Its not 9.5 lol

OpenStudy (anonymous):

K=4j-11 49=4j-11 +11 +11 ----------- 60=4j -- -- 4 4 j=15

OpenStudy (anonymous):

it is 15 lol

OpenStudy (anonymous):

ty